Find the solution to the IVP x4 y(4) – 4x®y +12x²y – 24xy +24y=0

Detailed answer
Answer #1

soli - Given ar y M_4n3y + 12 majl_24 any +24y=0 o ė Z. d Net Z=logn, then Then we have nyra 1-DY (DE off nry = D(D-1) y x31. z 47 :ysce the 27 e 27tz ettcue ega ce e²+265 + cs (e² ) ² + c (6 434 Garantcz. x² + 4x4 (rea. Joth) E ez =n; G, C2, C3, C